# Sampling Theory for Function Approximation with Numerical Redundancy

### v1

@@ 2.1 Numerical redundancy
We can conclude that if $\Psi$ is an overcomplete frame, the lower frame bound $A_{n}$ of the subsequence $\Phi_{n}$ will inevitably drop below $\operatorname{\epsilon_{\text{mach}}}^{2}\|\mathcal{T}_{n}\|_{2,H}=\operatorname{\epsilon_{\text{mach}}}^{2}B_{n}^{2}$ for sufficiently large $n$, resulting in a numerically redundant spanning set $\Phi_{n}$ according to Property 1. This is not the case for subsequences of Riesz bases [17, Theorem 3.6.6 (ii)].
2.2 Achievable accuracy of numerical approximations
As motivated in the previous section, when one computes the output of the operator $\mathcal{T}_{n}$ using a finite working precision $\operatorname{\epsilon_{\text{mach}}}$, the result can be modeled as the output of some other linear operator $\mathcal{X}_{n}$ that also maps from $\mathbb{C}^{n^{\prime}}$ to $H$ and that lies close to $\mathcal{T}_{n}$
@@ 2.2 Achievable accuracy of numerical approximations
For any $f\in H$, the error of the orthogonal projection onto the range of a linear operator $\mathcal{X}_{n}:\mathbb{C}^{n^{\prime}}\to H$ that satisfies (24) is bounded by
@@
Theorem 3 shows that one is guaranteed that accurate approximations are computable with finite precision, if an approximation in $V_{n}$ exists that is both accurate and can be expanded in $\Phi_{n}$ with reasonably sized coefficients $\mathbf{x}$. Furthermore, it shows that the analytical best approximation is computable up to an error of order $\operatorname{\epsilon_{\text{mach}}}$ if the frame bounds of $\Phi_{n}$ satisfy $A_{n}\approx B_{n}$. On the other hand, for numerically redundant $\Phi_{n}$, we have $\sqrt{B_{n}/A_{n}}\geq 1/\operatorname{\epsilon_{\text{mach}}}$ implying that accuracy cannot be guaranteed. In this case, the accuracy of the numerical approximation is generally significantly lower than the analytical best approximation in $V_{n}$.
Analogous to section 2.1, one can analyse what happens as $n\to\infty$ if the spanning sets $\Phi_{n}$ for the approximation spaces $\{V_{n}\}_{n\in\mathbb{N}}$ are subsequences of an infinite-dimensional set $\Psi=\{\psi_{i}\}_{i=1}^{\infty}$ as defined by (21). As previously discussed, if $\Psi$ is an overcomplete frame, the finite spanning sets $\Phi_{n}$ are numerically redundant as $n\to\infty$. According to Theorem 3, this means that high accuracy is not guaranteed for fixed $n$. However, the following theorem shows that convergence down to $\operatorname{\epsilon_{\text{mach}}}$ is guaranteed as $n\to\infty$. This is related to the existence of accurate approximations with reasonably sized expansion coefficients as $n\to\infty$.
@@
Using Theorem 3, (21) and $\lim_{n\to\infty}\|\mathcal{T}_{n}\|_{2,H}=\lim_{n\to\infty}\sqrt{B_{n}}=\sqrt{B}$ [5, Lemma 5], we find that
@@
Note that the infinite sum $\sum_{i=1}^{\infty}\psi_{i}x_{i}$ converges for all $\mathbf{x}\in\ell^{2}(\mathbb{N})$ since $\Psi$ is a Bessel sequence [17, Theorem 3.2.3]. Furthermore, it follows from (23) that there exist coefficients $\mathbf{c}\in\ell^{2}(\mathbb{N})$ satisfying $f=\sum_{i=1}^{\infty}\psi_{i}c_{i}$ with $\|\mathbf{c}\|_{2}\leq\|f\|_{H}/\sqrt{A}$. Hence, choosing $\mathbf{x}=\mathbf{c}$,
@@
2.3 Backward stability and the need for $\boldsymbol{\ell^{2}}$-regularization
In the previous sections, we introduced numerical redundancy and showed that rounding errors effectively change the approximation space. In this section, we analyse how to actually retrieve approximations, assuming we can make use of backward stable algorithms. In this setting, we will come to a similar yet stronger conclusion: accuracy is only guaranteed if the algorithm outputs expansion coefficients with reasonably sized coefficients. This naturally leads to an interest in $\ell^{2}$-regularized approximations. As mentioned in the introduction of this section, in a more practical setting this analysis would be performed after discretization.
The assumption of backward stability is common in the field of numerical analysis. It is mentioned in [34, section 1.5] that a method for computing $y=f(x)$ is called backward stable if, for any $x$, it produces a computed $\widehat{y}$ with a small backward error, that is, $\widehat{y}=f(x+\Delta x)$ for some small $\Delta x$. In what follows, we assume that small means that the perturbations can be bounded by $\|\Delta x_{i}\|\leq C\operatorname{\epsilon_{\text{mach}}}\|x_{i}\|$ for all inputs $x_{i}$, with $C>0$ a modest constant which may depend on the dimensions of $\{x_{i}\}_{i}$. In this case, we say that the algorithm is backward stable with stability constant $C$. Importantly, the perturbation $\Delta x$ does not only stem from rounding of the input $x$, but it also models the effect of rounding errors appearing throughout the calculation of $f(x)$.
First, we analyse the problem of computing the expansion coefficients of the orthogonal projection of a function $f$ onto $V_{n}$, i.e.,
@@ 2.3 Backward stability and the need for $\boldsymbol{\ell^{2}}$-regularization
Note that if $\Phi_{n}$ is linearly dependent, an infinite number of other sets of expansion coefficients exist for the orthogonal projection, yet the coefficients returned by (25) have the smallest $\ell^{2}$-norm. The following theorem shows that backward stability does not necessarily guarantee high accuracy for all $f\in H$ if $\Phi_{n}$ is numerically redundant. Its proof can be found in Appendix A.
@@
The previous theorem shows that backward stable algorithms for (25) give accurate results if $\Phi$ has a sufficiently large lower frame bound $A_{n}$. However, for numerically redundant spanning sets, no guarantees can be made. The error can be large, since $\widehat{\mathbf{c}}$ can grow unacceptably as $A_{n}$ goes to zero. Therefore, it is natural to force the solution to have a bounded $\ell^{2}$-norm by computing a regularized approximation instead
@@
where $\epsilon>0$ is a regularization parameter. Note that this mapping is again linear and that $\mathbf{c}$ is uniquely determined even if the spanning set $\Phi_{n}$ is linearly dependent. The following theorem shows that for sufficiently large $\epsilon$, backward stability does guarantee an error that is close to the achievable accuracy for numerical approximations, as identified in Theorem 3. The proof of Theorem 6 can be found in Appendix A.
@@
3 Discrete regularized least squares approximations

### v2

@@ 2.1 Definition and links to frames theory
We can conclude that if $\Psi$ is an overcomplete frame, the lower frame bound $A_{n}$ of the subsequence $\Phi_{n}$ will inevitably drop below $\operatorname{\epsilon_{\text{mach}}}^{2}\|\mathcal{T}_{n}\|_{2,H}=\operatorname{\epsilon_{\text{mach}}}^{2}B_{n}^{2}$ for sufficiently large $n$, resulting in a numerically redundant spanning set $\Phi_{n}$ according to Property 1. This is not the case for subsequences of Riesz bases [19, Theorem 3.6.6 (ii)].
2.2 Best approximation in finite-precision arithmetic
As motivated in the previous section, when working with a finite working precision $\operatorname{\epsilon_{\text{mach}}}$, the output of the operator $\mathcal{T}_{n}$ can be modeled as the output of some other linear operator $\mathcal{X}_{n}$ that also maps from $\mathbb{C}^{n^{\prime}}$ to $H$ and that lies close to $\mathcal{T}_{n}$
@@ 2.2 Best approximation in finite-precision arithmetic
For any $f\in H$, the best approximation error in the range of a linear operator $\mathcal{X}_{n}:\mathbb{C}^{n^{\prime}}\to H$ that satisfies (24) is bounded by
@@
Theorem 3 shows that one is guaranteed that accurate approximations are computable in the presence of rounding errors, if an approximation in $V_{n}$ exists that is both accurate and can be expanded in $\Phi_{n}$ with reasonably sized coefficients $\mathbf{x}$. Furthermore, it shows that the analytical best approximation is computable up to an error of order $\operatorname{\epsilon_{\text{mach}}}$ if the frame bounds of $\Phi_{n}$ satisfy $A_{n}\approx B_{n}$. On the other hand, for numerically redundant $\Phi_{n}$, we have $\sqrt{B_{n}/A_{n}}\geq 1/\operatorname{\epsilon_{\text{mach}}}$ such that accuracy is not guaranteed. Of course, this is only an upper bound on the error. The following lemma showcases that the bound is indeed relevant. In particular, for numerically redundant sets the upper bound can always be attained.
Lemma 1.
If $\Phi_{n}$ is numerically redundant, there exist a linear operator $\mathcal{X}_{n}:\mathbb{C}^{n^{\prime}}\to H$ that satisfies (24), a nonzero function $f\in H$, and nonzero coefficients $\mathbf{c}\in\mathbb{C}^{n^{\prime}}$ such that
$\|f-\widehat{P}_{n}f\|_{H}=\|f-\mathcal{T}_{n}\mathbf{c}\|_{H}+\sigma\|\mathbf{c}\|_{2},$
where $\sigma$ is the largest singular value of $\mathcal{T}_{n}$ that is smaller than or equal to $\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}$, implying $0<\sigma\leq\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}$.
Proof.
The operator $\mathcal{T}_{n}$ admits a singular value decomposition, i.e.,
$\mathcal{T}_{n}=\sum_{i=1}^{n}\sigma_{i}u_{i}\mathbf{v}_{i}^{*}$
where $\{u_{i}\}_{i}$ is an orthonormal basis for $V_{n}=\operatorname{R}(\mathcal{T}_{n})$ and $\{\mathbf{v}_{i}\}_{i}$ is an orthonormal basis for $\mathbb{C}^{n^{\prime}}$. Denote by $r$ the number of singular values of $\mathcal{T}_{n}$ that are strictly greater than $\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}$. For numerically redundant $\Phi_{n}$, one has $r<n$. Consider now
$\mathcal{X}_{n}=\sum_{i=1}^{r}\sigma_{i}u_{i}\mathbf{v}_{i}^{*}$
and $f=\mathcal{T}_{n}\mathbf{c}$ with $\mathbf{c}=C\mathbf{v}_{r+1}$ for some nonzero $C$. One can check that (24) holds for this choice of $\mathcal{X}_{n}$. Furthermore,
$\|f-\widehat{P}_{n}f\|_{H}=\|f\|_{H}=\|\mathcal{T}_{n}\mathbf{c}\|_{H}=C\sigma_{r+1}\|u_{r+1}\|_{H}=\|f-\mathcal{T}_{n}\mathbf{c}\|_{H}+C\sigma_{r+1},$
where in the last step we add $\|f-\mathcal{T}_{n}\mathbf{c}\|_{H}=0$. ∎
Analogous to section 2.1, one can analyse what happens as $n\to\infty$ if the spanning sets $\Phi_{n}$ for the approximation spaces $\{V_{n}\}_{n\in\mathbb{N}}$ are subsequences of an infinite-dimensional set $\Psi=\{\psi_{i}\}_{i=1}^{\infty}$ as defined by (21). If $\Psi$ is an overcomplete frame, recall that the finite spanning sets $\Phi_{n}$ are numerically redundant for sufficiently large $n$. This means that some functions in the approximation space cannot be represented with small expansion coefficients and, hence, it follows from Theorem 3 that accuracy of numerical approximations is not guaranteed. On the other hand, the following theorem shows that convergence down to $\operatorname{\epsilon_{\text{mach}}}$s guaranteed as $n\to\infty$. This is related to the existence of accurate approximations with reasonably sized expansion coefficients in the set $\Psi$. This is an important result: even though the spanning sets $\Phi_{n}$ are numerically redundant for finite $n$, we can still guarantee convergence of numerical approximations down to $\operatorname{\epsilon_{\text{mach}}}$ for every $f\in H$ using the frame property of the infinite-dimensional set.
@@
Using Theorem 3, (21) and $\lim_{n\to\infty}\|\mathcal{T}_{n}\|_{2,H}=\lim_{n\to\infty}\sqrt{B_{n}}=\sqrt{B}$ [6, Lemma 5], we find that
@@
Note that the infinite sum $\sum_{i=1}^{\infty}\psi_{i}x_{i}$ converges for all $\mathbf{x}\in\ell^{2}(\mathbb{N})$ since $\Psi$ is a Bessel sequence [19, Theorem 3.2.3]. Furthermore, it follows from (23) that there exist coefficients $\mathbf{c}\in\ell^{2}(\mathbb{N})$ satisfying $f=\sum_{i=1}^{\infty}\psi_{i}c_{i}$ with $\|\mathbf{c}\|_{2}\leq\|f\|_{H}/\sqrt{A}$. Hence, choosing $\mathbf{x}=\mathbf{c}$,
@@
3 Interpreting numerical approximations as $\boldsymbol{\ell^{2}}$-regularized approximations
In this section, we motivate why it is natural to model numerical approximations as $\ell^{2}$-regularized approximations.
3.1 Error analysis with backward stability assumption
What can one reasonably expect from the output of a numerical algorithm? Due to rounding errors, recovering the exact solution is generally not possible. Moreover, even obtaining an output close to the exact solution is often unrealistic: small rounding errors can be significantly amplified by the ill-conditioning of a problem—a property that is independent of the algorithm. In numerical analysis, it is therefore common to assume / require that a numerical method is backward stable. More specifically, a method for computing $y=g(x)$ is called backward stable if, for any $x$, it produces a computed $\widehat{y}$ with a small backward error, that is, $\widehat{y}=g(x+\Delta x)$ for some small $\Delta x$ [35, section 1.5]. Here, the perturbation $\Delta x$ does not only model rounding of the input $x$, but it also models the effect of rounding errors appearing throughout the calculation of $g(x)$. It is now natural to wonder whether backward stable algorithms output satisfying solutions in our setting. In what follows, we assume that small means that the perturbations can be bounded by $\|\Delta x_{i}\|\leq C\operatorname{\epsilon_{\text{mach}}}\|x_{i}\|$ for all inputs $x_{i}$, with $C>0$ a modest constant which may depend on the dimensions of $\{x_{i}\}_{i}$. In this case, we say that the algorithm is backward stable with stability constant $C$.
In essence, we are interested in computing the expansion coefficients of the orthogonal projection of a function $f$ onto $V_{n}$, i.e.,
@@ 3.1 Error analysis with backward stability assumption
Note that if $\Phi_{n}$ is linearly dependent, an infinite number of other sets of expansion coefficients exist for the orthogonal projection, yet the coefficients returned by (25) have the smallest $\ell^{2}$-norm. The following theorem shows what accuracy can be expected from the output of a backward stable algorithm that implements the mapping (25). Its proof can be found in Appendix A.
@@
The theorem shows that if approximations are computed using a spanning set with $A_{n}\approx B_{n}$, numerical rounding errors are negligble, i.e., of order $\operatorname{\epsilon_{\text{mach}}}$. However, for spanning sets with $A_{n}\ll B_{n}$ this is not the case. More specifically, accuracy can only be guaranteed if there exists a set of coefficients $\mathbf{x}$ such that $\|f-\mathcal{T}_{n}\mathbf{x}\|_{H}+C\operatorname{\epsilon_{\text{mach}}}\|\mathcal{T}_{n}\|_{2,H}\|\mathbf{x}\|_{2}$ remains small, and when $C\operatorname{\epsilon_{\text{mach}}}\|\widehat{\mathbf{c}}\|_{2}$ is sufficiently small where $\widehat{\mathbf{c}}$ is the computed solution. Following the discussion in section 2.2, the first condition is natural when working in finite precision, independently of which method is used to compute coefficients. On the other hand, the second condition on $\|\widehat{\mathbf{c}}\|_{2}$ is new and non-trivial. To see why, note that the coefficients of the best approximation $\mathbf{c}=\mathcal{T}_{n}^{\dagger}f$ can grow arbitrarily large, in particular larger than $\|f\|_{H}/(\|\mathcal{T}_{n}\|_{H}\operatorname{\epsilon_{\text{mach}}})$ for numerically redundant spanning sets. Consequently, the norm of the computed coefficients $\widehat{\mathbf{c}}$ will likely be large as well and no accuracy can be guaranteed.
Does this mean that we cannot obtain any guarantees on the accuracy of numerical approximations when using a numerically redundant spanning set? The answer is no. However, we must slightly adapt our problem statement to reflect what is actually computed when standard solvers, such as MATLAB’s backslash, are applied to numerically rank-deficient system. Namely, in such cases, $\ell^{2}$-regularization is used—even if the user is not explicitly aware of it. We therefore consider the regularized formulation:
@@
where $\epsilon>0$ is a regularization parameter. The mapping remains linear, and $\mathbf{c}$ is now uniquely determined even if the spanning set $\Phi_{n}$ is linearly dependent. It is important to emphasize that our use of $\ell^{2}$-regularization differs from more typical settings, e.g. in inverse problems, where it is introduced as a modeling choice to encode prior knowledge such as smoothness of the solution. In contrast, in our setting regularization is inevitable in order to ensure numerical computability. In this sense, it is not a matter of choice but a numerical necessity.
The following theorem states accuracy guarantees for backward stable algorithms, now implementing the mapping (26). The proof of Theorem 6 can again be found in Appendix A.
@@
We are now guaranteed to obtain accurate numerical approximations provided that there exists a set of coefficients $\mathbf{x}$ such that $\|f-\mathcal{T}_{n}\mathbf{x}\|_{H}+\epsilon\|\mathbf{x}\|_{2}$ is sufficiently small, where $\epsilon$ is on the order of $\operatorname{\epsilon_{\text{mach}}}$. This condition is natural in finite-precision arithmetic, as demonstrated in section 2.2.
3.2 Implicit regularization in numerical orthogonalization
From the discussion above, it may be tempting to always try to avoid numerical redundancy and obtain an orthonormal basis for $V_{n}$. Indeed, several computational methods operate with a numerically orthogonalized basis, see for example [42, 25]. But do these methods produce a substantially different approximation from that given by (26)? In this section, we show that regularization is also implicitly present in straightforward numerical orthogonalization techniques. Our goal is to convey the main ideas, though each of the pointers provided deserves further exploration. Finally, we note that Theorem 3 also suggests it is fundamentally impossible to numerically orthogonalize a redundant set without encountering issues related to finite-precision arithmetic.
For simplicity, we assume that a set of linearly independent functions $\{\eta_{1},\dots,\eta_{n}\}$ is given, which span the approximation space $V_{n}$ with $\dim(V_{n})=n$. One straightforward way of orthogonalizing consists of discretizing these functions on a fine grid $\{\mathbf{t}_{j}\}_{j=1}^{l}$ and computing a factorization
This basis is orthogonal with respect to the discrete inner product defined by
$\langle f,g\rangle_{l}=\frac{1}{l}\sum_{j=1}^{l}f(\mathbf{t}_{j})\overline{g(\mathbf{t}_{j})}.$
If the grid $\{\mathbf{t}_{j}\}_{j=1}^{l}$ is sufficiently dense, the functions are also (approximately) orthonormal with respect to the inner product associated with the Hilbert space $H$. The basis $\{\widetilde{\eta}_{j}\}_{j=1}^{n}$ can then be used as a spanning set $\Phi_{n}=\{\phi_{i,n}\}_{i=1}^{n}$ for least squares approximation.
Importantly, in the case of a numerically redundant set of functions $\{\eta_{1},\dots,\eta_{n}\}$, the matrices $A$ and $R$ are heavily ill-conditioned. More specifically, their singular value profiles mimick that of the synthesis operator associated with $\{\eta_{1},\dots,\eta_{n}\}$, having very small yet nonzero singular values, as discussed in section 2. Can (28) even be computed with high accuracy in such cases? As an example, we consider the QR decomposition as a factorization method for (27), such that $R$ is upper-triangular. The accuracy of solving triangular systems of equations is analysed extensively in [34], where it is shown that the forward error depends on the Skeel condition number [49]:
[…]
In this case, we do not have to solve an ill-conditioned linear system; however, we can only access the basis $\{\widetilde{\eta}_{1},\dots,\widetilde{\eta}_{n}\}$ on the grid $\{\mathbf{t}_{j}\}_{j=1}^{l}$. As an example, we consider the accuracy of the very standard and backward stable Householder QR algorithm, which is analysed in [35, section 19.3]. Following [35, p.361], one is guaranteed to compute matrices $\hat{Q}$ and $\hat{R}$ such that
where $C$ is a constant that depends on the dimensions of the problem and the working precision $\operatorname{\epsilon_{\text{mach}}}$ and $a_{j}$ denotes the $j$-th column of $A$. Due to the small singular values of $A$, the range of $A+\Delta A$, and consequently the range of $\hat{Q}$, can differ significantly from that of $A$. A toy example illustrating this phenomenon is given in [51, p. 263], and a detailed study of the angle between the subspaces $\operatorname{R}(A)$ and $\operatorname{R}(A+\Delta A)$ is given in [27]. In conclusion, there are no guarantees that the computed basis functions (the columns of $\hat{Q}$) span the full approximation space $V_{n}$ (the range of $A$).
Remark 5.
Observe the similarity between (30) and a truncated QR decomposition—a form of $\ell^{2}$-regularization: the range of $\hat{Q}$ contains the singular vectors of $A$ associated with large singular values ($\sigma_{i}\gg\operatorname{\epsilon_{\text{mach}}}$), while the directions associated with small singular values may be significantly perturbed or entirely missing. As a result, the best approximation cannot, in general, be computed.
Finally, we note that there exist numerical algorithms that do succeed in computing an orthonormal basis for the whole approximation space $V_{n}$. However, in order to do so they do not explicitly orthogonalize a numerically redundant set. Instead, they rely on additional analytical knowledge. One example is the Stieltjes procedure for generating orthogonal polynomials [28], which recently gained popularity for numerical polynomial approximations in the form of the Vandermonde with Arnoldi (VwA) algorithm [18]. The orthogonalization procedure is summarized in Algorithm 1.
Algorithm 1 Stieltjes procedure for generating orthonormal polynomials.
$\eta_{1}\leftarrow 1$, $\widetilde{\eta}_{1}\leftarrow\eta_{1}/\|\eta_{1}\|_{H}$
for $i=2$ to $n$ do
     $\eta_{i}(\mathbf{x})\leftarrow\mathbf{x}\;\widetilde{\eta}_{i-1}(\mathbf{x})$
     $v\leftarrow\eta_{i}-\sum_{j=1}^{i-1}\langle\eta_{i},\widetilde{\eta}_{j}\rangle_{H}\;\widetilde{\eta}_{j}$
     $\widetilde{\eta}_{i}\leftarrow v/\|v\|_{H}$
end for
return $\{\widetilde{\eta}_{1},\dots,\widetilde{\eta}_{n}\}$
The success of this procedure stems from the fact that one does not orthogonalize monomials (which are numerically redundant for sufficiently large $n$ as demonstrated on Figure 6(a) and 6(b)) yet iteratively builds “Arnoldi” basis functions $\eta_{i}$ that are more suitable for orthogonalization. Indeed, Figure 1 shows that the set $\eta_{i}$ defined by Algorithm 1 for the domain $[-1,1]$ has frame bounds satisfying $A_{n}\approx B_{n}$. Hence, the procedure does not suffer from the effect of finite-precision arithmetic. Stability of orthogonalization of the Arnoldi basis is analysed for a variety of domains in [53, 52]. We remark that the VwA algorithm performs discrete orthogonalization instead of the continuous orthogonalization in Algorithm 1. Furthermore, note that the algorithm can be extended to other sets of functions provided they have similar Vandermonde-like structure, as illustrated in section 3.3.
We have shown that obtaining an orthonormal basis for the full approximation space is non-trivial and can only be done numerically in specific cases, such as with the VwA algorithm. Even when an algorithm is available that avoids the effects of finite-precision arithmetic, it is natural to ask about the advantages and disadvantages of using an orthogonalized spanning set versus a regularized approximation with a redundant spanning set. Most importantly, as illustrated in section 3.3, regularization typically slows down convergence. On the other hand, it reduces the number of measurements needed for accurate approximation. To fairly compare convergence behaviour as a function of the number of samples, this phenomenon must be examined more closely, which is exactly the focus of the current paper. Additionally, orthogonalization often throws away exploitable structure. For example, in the case discussed below, the discrete least squares matrix associated with $\Phi_{n}$ is a submatrix of a Fourier matrix. This structure can be leveraged to design efficient least squares methods—see, for instance, the AZ algorithm for frame approximations [22, 31]—or to enable cheaper evaluation of the approximation.
(a) first eight basis functions
(b) singular value profile of $\mathcal{T}_{n}$ for $n=100$
Figure 1: The Arnoldi basis associated with Stieltjes orthogonalization / the Vandermonde with Arnoldi algorithm along with the singular value profile of the associated synthesis operator. This set of functions has well behaved frame bounds, i.e., $A_{n}\approx B_{n}$, in contrast to the monomials displayed in Figure 6(a) and 6(b). Hence, the set can be numerically orthogonalized without suffering from the effects of finite precision.
3.3 Numerical evidence
In section 3.1, we showed that while practical algorithms cannot reliably compute the best approximation, they can compute $\ell^{2}$-regularized approximations. Furthermore, section 3.2 demonstrated that $\ell^{2}$-regularization cannot be avoided using standard numerical orthogonalization techniques. In this section, we illustrate these theoretical findings with a numerical example.
To this end, we reproduce [18, Example 3]. More specifically, we approximate the function $f(x)=1/(10-9x)$ on $[-1,1]$ using a Fourier extension on $[-2,2]$, i.e.,
$f(x)\approx\sum_{k=-n}^{n}c_{k}\exp(ik\pi x/2).$
For more information on Fourier extensions, see §6.2. For now, note that this is a numerically redundant spanning set. Since $f$ is real, this can be simplified to
On [18, Figure 5.1], a comparison is shown between the direct computation of a least squares fit using MATLAB’s backslash and the Vandermonde with Arnoldi (VwA) algorithm. The computations are based on function evaluations in 1000 Chebyshev points. Note that VwA is applicable in this setting, since the discrete least squares matrix has Vandermonde structure as described in [18, Example 3].
We add some extra methods to the comparison: approximation using a numerically orthogonalized basis obtained via a QR decomposition as described above, and a truncated singular value (TSVD) approximation, which is an explicit way of computing an $\ell^{2}$-regularized approximation. For implementational details, we refer to [18] and [32]. The results are presented in Figure 2. The VwA algorithm achieves exponential convergence, which is to be expected in $\operatorname{span}(\Phi_{n})$ [36]. On the other hand, the accuracy of the approximation computed by MATLAB’s backslash, as well as of the regularized and numerically orthogonalized approximations, begins to deviate from that of the best approximation starting at $n=20$. This example confirms that standard routines such as MATLAB’s backslash and straightforward numerical orthogonalization techniques implicitly regularize the approximation problem. Furthermore, as shown in Figure 3, the spanning set $\Phi_{n}$ becomes numerically redundant starting at $n=20$. Hence, this illustrates that the concept of numerical redundancy accurately describes when regularization significantly impacts accuracy.
4 Error analysis of regularized approximations